Freon – 12, CCl2F2, is prepared from CCl4 by reaction with HF. The other product of this reaction is HCl. Outline the steps needed to determine the percent yield of a reaction that produces 12.5 g of CCl2F2 from 32.9 g of CCl4. Freon – 12 has been banned and is no longer used as a refrigerant because it catalyzes the decomposition of ozone and has a very long lifetime in the atmosphere. Determine the percent yield.
1
Expert's answer
2020-11-22T11:12:38-0500

CCl4 + 2 HF "\rightarrow" CCl2F2 + 2 HCl

1mole of CCl4 is responsible for production of 1 mole of CCl2F2

Moles of CCl4 = "\frac{32.9}{153.8}" = 0.254 moles


Theoritical yield of CCl2F2 formed = 0.254× 120.9 = 30.81g


%yield = "\frac{12.5×100}{30.81}"

= 40.57%